OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9(F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 25 degrees Celsius is.

OpenStudy (johnweldon1993):

so just rearrange this so f = something C = 5/9(F-32) first multiply both sides by 9/5 (the reciprocal o 5/9 thus canceling this out) 9/5C = F - 32 add 32 to both sides to make F be alone F = 9/5C + 32 if this is true...then Part 2: Determine how many degrees Fahrenheit 25 degrees Celsius is. we take our equation F = 9/5C + 32 and plug in 25 for C so F = 9/5(25) + 32 F = 225/5 + 32 F = 45 + 32 F = 77